### v1

# To reset or not to reset in a finite domain: that is the question

Abstract
We investigate the search of a target with a given spatial distribution in a finite one-dimensional domain. The searcher follows Brownian dynamics and is always reset to its initial position when reaching the boundaries of the domain (boundary resetting). In addition, the searcher may be reset to its initial position from any internal point of the domain (bulk resetting). Specifically, we look for the optimal strategy for bulk resetting, i.e., the spatially dependent bulk resetting rate that minimizes the average search time. The best search strategy exhibits a second-order transition from vanishing to non-vanishing bulk resetting when varying the target distribution. The obtained mathematical criteria are further analyzed for a monoparametric family of distributions, to shed light on the properties that control the optimal strategy for bulk resetting. Our work paves new research lines in the study of search processes, emphasizing the relevance of the target distribution for the optimal search strategy, and identifies a successful framework to address these questions.
Nature is full of instances of first-passage processes [1, 2]. Search problems can be formulated in terms of them, focusing on analyzing specific quantities related to the first-passage time—like the mean first-passage time (MFPT)—as key indicators of search performance. These problems are especially relevant from an applied perspective and have a significant impact across a wide range of disciplines, from biophysics—at various scales, from single molecules [3] to animal foraging [4, 5]—to computing methods [6], rescue mission strategies [7], or reliability theory [8, 9], to name just a few.
The typical goal in a search problem is the optimization of the search strategy, e.g., finding out the strategy that minimizes the MFPT [10]. Intermittent strategies that combine reactive random search phases with faster, non-reactive relocation phases, have proven to be highly effective to this end [11]. In its simplest introduction, stochastic resetting constitutes a specific instance of an intermittent strategy, with instantaneous relocations [12, 13]. In the last years, optimization of stochastic resetting has been a prolific line of research [14, 13, 15, 16, 17, 18, 9, 19, 20, 21, 22, 23, 24]. The original simple proposal for stochastic resetting has been sophisticated by implementing more realistic reset mechanisms, e.g., using confining potentials [25, 18, 26, 27, 20, 28] or introducing some cost for the resets [29, 30, 9, 31, 32, 33].
Remarkably, most of optimization problems within search research have been posed for a target with a well-defined position, with very few exceptions [10, 13, 23]. Yet, optimizing the search strategy for a fixed target position is basically pointless for a practical search problem: the optimal strategy thus found depends on the target position, of which an actual searcher is uncertain. Therefore, the role of the degree of uncertainty about the target location—codified in its spatial distribution—is a significant issue that demands to be investigated to develop optimal strategies in more realistic contexts. Moreover, it must be remarked that stochastic resetting has been predominantly studied in infinite domains, where resets are always beneficial because they “kill” the trajectories drifting away from the target. For finite domains [34, 35, 36, 9, 37], resetting has been observed to be detrimental for some specific situations.
In this Letter, we address the gaps described above by analyzing the impact on the optimal resetting strategy of the target’s spatial distribution in a finite domain. To this end, we employ the paradigmatic model of a resetting Brownian searcher moving in a finite one-dimensional box. In this context, we will look for the best resetting strategy, allowing for it to be inhomogeneous, i.e., the resetting rate depends in general on the position of the searcher. Space-dependent resetting has been considered in systems with a localized target [14], and the stationary state obtained by using a quantum mechanical mapping [38].
For the considered one-dimensional search, a bounded distribution naturally leads to the concept of resetting boundaries: if the searcher reaches either wall of the box without finding the target, the optimal strategy is clearly to restart the search from its initial position. Finally, each target distribution determines an optimal—potentially inhomogeneous—bulk resetting rate. Here, we show how this optimal rate can be derived from the target distribution, and we provide clear criteria for when bulk resetting can be avoided.
Model.- We consider a Brownian searcher that freely diffuses in a one-dimensional box. Let $X(t)$ be its position at time $t$, which is looking for the target located at $X_{T}$. From a mathematical point of view, $X$ is a stochastic process and $X_{T}$ is a random variable with a certain distribution $P$. The distribution $P$ quantifies the degree of uncertainty about the location of the target; flat and single delta-peaked distributions, respectively, represent extreme cases of full uncertainty and full certainty. The target remains motionless in any trajectory, which has been referred to as a quenched-disorder target [23] due to its similarity to disordered systems  [^1] As opposed to an annealed-disorder target, whose position would evolve in time..
The searcher starts from the center of the box, $x=0$. When it reaches the target at $x=x_{T}$, the dynamics terminates. Until then, its dynamics—illustrated in Fig. 1—is as follows: (i) When hitting the boundaries at $x=\pm\ell$, the position is instantaneously reset to the center; (ii) the probability per unit time $r$ of a reset event in the bulk depends solely on space, i.e., $r\equiv r(x)$, and the reset event instantaneously restarts the search from the center; (iii) between reset events, the searcher freely diffuses with a constant diffusion coefficient $D$.
Figure 1: Sketch of a single trajectory for the Brownian search of a spatially distributed target. Left: One trajectory starting from $x=0$. The blue solid trace stands for diffusive motion, with instances of a boundary reset (vertical green stroke) and bulk resets (vertical yellow strokes). The target position $x_{T}$ (horizontal dotted red line) has been drawn from its distribution $P(x_{T})$. Domain boundaries are at $\pm\ell$ (horizontal dashed green lines). Right: Spatial distribution of the target.
On the one hand, our resetting boundaries clearly expedite—with respect to reflecting ones—the search in a one-dimensional domain: reaching one wall without success means that the target lies on the other half of the box. On the other hand, bulk resets at a certain position $x$ are beneficial (detrimental) if $x\,x_{T}$ is negative (positive). Therefore, two natural questions arise: When is bulk resetting advantageous for finding the target, given its distribution $P(x_{T})$? What is the optimal bulk resetting strategy $r(x)$ for a given $P(x_{T})$? The main goal of this Letter is to answer these two key questions.
For search optimization, we start by studying the MFPT $\tau(x_{0}|x_{T})$ from an arbitrary initial position $x_{0}$ to the target at a fixed location $x_{T}$. Specifically, we focus on the MFPT $\tau_{0}(x_{T})\equiv\tau(0|x_{T})$ from the center of the box to the target. Then, we consider its average with respect to the target distribution
[…]
where we have introduced the notation $\langle\cdot\rangle_{T}$ for averages with respect to $P(x_{T})$. In this Letter, the average time $\overline{\tau}[r]$ is the main figure of merit for quantifying the success of the search. Note that $\overline{\tau}[r]$ is a functional of the bulk resetting strategy $r(x)$ through $\tau_{0}$—the functional dependence of the latter on $r(x)$ has not been explicitly included in the notation to avoid cluttering the formulas.
The MFPT at fixed $x_{T}$ can be obtained by solving the backward Fokker-Planck equation
$\partial_{x_{0}}^{2}\tau(x_{0}|x_{T})$ $=-1+r(x_{0})[\tau(x_{0}|x_{T})-\tau(0|x_{T})],$ (2a)
$\tau(\pm 1|x_{T})$ $=\tau(0|x_{T}),\qquad\tau(x_{T}|x_{T})=0,$ (2b)
where we have switched to dimensionless variables, respectively taking $\ell$ and $\ell^{2}/D$ as units for space and time, as detailed in Sec. I of [40]. The boundary conditions (2b) implement resetting boundaries at the walls and the absorbing point at the target position, respectively. Since we are interested in $\tau_{0}(x_{T})$, it suffices to solve Eq. (2) for $x_{0}$ in the interval between one of the walls and the target position that includes $x=0$. While the analytical solution for any bulk resetting rate $r(x)$ is out of reach, it can still be obtained for specific choices of $r(x)$.
Homogeneous resetting.- Here, we consider the average MFPT defined in Eq. (1) for a homogeneous bulk resetting rate $r$. The average MFPT is then no longer a functional, but a standard function of $r$—we thus write $\overline{\tau}(r)$. We will optimize $\overline{\tau}(r)$ for a given $P$, i.e., we look for the best homogeneous resetting rate $r^{*}$ that minimizes $\overline{\tau}$. As already stated above, the dependence on $r$ enters through the solution for $\tau_{0}$ from Eq. (2), which can be solved for constant $r$—see Sec. II of [40]:
$\tau_{0}(x_{T};r)=r^{-1}\left[\frac{\cosh[(|x_{T}|+\frac{1}{2})\sqrt{r}]}{\cosh[\sqrt{r}/2]}-1\right].$ (3)
To obtain $\overline{\tau}(r)$, it suffices to compute the average of Eq. (3) with respect to $P(x_{T})$. We note that $\tau_{0}(x_{T})$ is a convex function that diverges for $r\to\infty$. Since $r\geq 0$, the behavior of $\overline{\tau}(r)$ close to $r=0$ depends on the sign of $d\overline{\tau}/dr|_{r=0}$. For positive derivative, $\overline{\tau}(r)$ has a local minimum at $r=0$, so the best search strategy implies no resetting in the bulk, $r^{*}=0$. Instead, for negative derivative, $\overline{\tau}(r)$ has a local maximum at $r=0$ and there appears an optimum non-zero resetting rate, $r^{*}\neq 0$.
The above discussion implies that the key quantity is
[…]
Specifically, $m>0$ ($m<0$) entails that the minimum of $\overline{\tau}$ is attained at $r^{*}=0$ ($r^{*}>0$). Using Eq. (3), we get
$M(x_{T})=\frac{1}{24}\left[x_{T}^{4}-\absolutevalue{x_{T}}\left(1-2x_{T}^{2}\right)\right].$ (5)
As shown in the top panel of Fig. 2, $M(x_{T})<0$ for $|x|<x_{T}^{c}\equiv(\sqrt{5}-1)/2$, whereas $M(x_{T})>0$ for $x_{T}^{c}<|x|<1$. Therefore, “small”, close to the center, values of $x_{T}$ contribute to destabilize the no-resetting strategy in the bulk, whereas “large”, close to the walls, values of $x_{T}$ contribute to stabilize it.
Figure 2: Derivative controlling the stability of the no-resetting strategy. Positive (negative) values, represented as warm (cold) colors, thereof contribute to zero (nonzero) bulk resetting as the best search strategy. Top: Standard derivative $M(x_{T})$ defined in Eq. (5) for the homogeneous case. The change of stability is reached at $x_{T}^{c}$ (black diamond). Bottom: Symmetrized functional derivative $M_{s}(x_{T},x)$ defined in Eq. (To reset or not to reset in a finite domain: that is the question) for the inhomogeneous case. White color indicates saturated values. The contour line at which $M_{s}=0$ is marked (dashed line), as well as the value $x_{T}^{(0)}$ (black diamond) below which $M_{s}(x_{T},x)<0$, $\forall x$.
It is worth noting some general results. First, for a flat distribution, $m>0$: in a finite domain, homogeneous resetting does not give any advantage for a homogeneously distributed target. Second, $m<0$ for any target distribution with finite support in the subinterval $0\leq x_{T}<x_{T}^{c}$, because $M(x_{T})<0$ where $P(x_{T})\neq 0$: therein, homogeneous resetting in the bulk is advantageous and $r^{*}\neq 0$. Conversely, $m>0$ for any target distribution with finite support in the subinterval $x_{T}^{c}<x_{T}<1$, because $M(x_{T})>0$ where $P(x_{T})\neq 0$: therein, homogeneous resetting in the bulk is detrimental and $r^{*}=0$.
To further illustrate the discussion above, we now introduce a specific monoparametric family of symmetric target distributions. In particular, we choose a variant of the beta distribution,
$P(x_{T};\beta)=\frac{\Gamma\left(\beta+\frac{1}{2}\right)}{\sqrt{\pi}\Gamma(\beta)}(1-x_{T}^{2})^{\beta-1},$ (6)
defined for $-1<x_{T}<1$, with $\beta>0$. This $\beta$-family interpolates between a single delta peak at the center of the box for $\beta\to+\infty$ and two equally weighted delta peaks at the boundaries for $\beta\to 0$, passing through the flat distribution for $\beta=1$. Introducing Eqs. (3) and (6) into Eq. (1), it is possible to analytically obtain the average MFPT, as detailed in Sec. II of [40]. Yet, for our discussion, it suffices to obtain the linear coefficient $m$ in Eq. (4), which is now a function of $\beta$:
$m(\beta)=\frac{1}{8[3+4\beta(\beta+2)]}-\frac{(\beta-1)\Gamma(\beta+\frac{1}{2})}{24\sqrt{\pi}\Gamma(\beta+2)}.$ (7)
We note that $m(\beta)$ only vanishes at $\beta=\beta_{c}\simeq 1.71$. For $\beta<\beta_{c}$, we have $m>0$ and $\overline{\tau}(r)$ is minimum at $r=0$. Instead, for $\beta>\beta_{c}$, we have $m<0$ and $\overline{\tau}(r)$ has a local maximum at $r=0$. Therefore, the global minimum of $\overline{\tau}(r)$ is attained at a certain value $r^{*}>0$. This behavior is reminiscent of a continuous phase transition, with the optimal homogeneous resetting rate $r^{*}$ playing the role of the order parameter: $r^{*}=0$ for $\beta<\beta_{c}$, $r^{*}\neq 0$ for $\beta>\beta_{c}$, with $r^{*}=0$ for $\beta=\beta_{c}$.
For the homogeneous case, we have proven that nonresetting in the bulk is the optimal strategy for certain target distributions, specifically those for which the region with positive $M(x_{T})$ (close to the walls) is more important, when weighted with the target distribution $P(x_{T})$, than the region with negative $M(x_{T})$ (close to the center). For the monoparametric family (6), this corresponds to $\beta<\beta_{c}$: loosely speaking, the weight of targets close to the center is insufficient to destabilize the strategy with $r=0$ in the bulk. Indeed, this is physically sensible: since the resetting boundaries already facilitate the search, extra resetting in the bulk is beneficial only if the weight of targets close to the center is large enough.
Optimization of heterogeneous resetting.- We now move to the general case with heterogeneous resetting $r(x)$. Similarly to the homogeneous case, whether or not bulk resetting is advantageous is controlled by the derivative of the average MFPT with respect to the resetting rate, evaluated at $r(x)\equiv 0$. However, now this quantity is a functional derivative, specifically
$\mu(x)$ $\equiv\left.\frac{\delta\overline{\tau}}{\delta r(x)}\right|_{r(x)\equiv 0}=\expectationvalue{M(x_{T},x)}_{T},$ (8)
$M(x_{T},x)$ $\equiv\left.\frac{\delta\tau_{0}(x_{T})}{\delta r(x)}\right|_{r(x)\equiv 0},$ (9)
instead of the standard derivative in Eq. (4). This can be computed analytically, as detailed in Sec. III of [40]:
$M(x_{T},x)=$ $\frac{x_{T}x\left(1-|x|\right)^{2}}{2}\Theta\left(-x_{T}x\right)$
$+\frac{x\left(x_{T}-x\right)\left(1+|x|\right)}{2}\Theta\left(x_{T}x\right)\Theta\left(|x_{T}|-|x|\right)\!\!,$ (10)
where $\Theta$ stands for Heaviside’s step function. The two terms on the right hand side of Eq. (To reset or not to reset in a finite domain: that is the question) have opposite signs: the first is always negative, whereas the second is always positive. Bulk resetting at $x$ is thus beneficial for $x_{T}x<0$ (target and particle positions at opposite sides of the box), whereas it is detrimental for $x_{T}x>0$ and $|x_{T}|>|x|$ (particle between the center and the target). For $x_{T}x>0$ and $|x_{T}|<|x|$, $M(x_{T},x)$ vanishes.
The sign of $\mu(x)$ provides a transparent criterion to elucidate the optimality of the no-resetting strategy in the bulk. It is guaranteed that $\overline{\tau}$ attains a local minimum for $r(x)=0$ if $\mu(x)>0$ for all $x$. Otherwise, the optimal search strategy involves nonvanishing resetting in the bulk. We will now focus on symmetric target distributions, i.e., $P(x_{T})=P(-x_{T})$. Therein, the optimal resetting strategy inherits the symmetry of the problem and it suffices to analyze the symmetrized function $M_{s}(x_{T},x)\equiv[M(x_{T},x)+M(-x_{T},x)]/2$ and its average with respect to $P(x_{T})$ in the interval $[0,1]$.
The symmetrized function $M_{s}(x_{T},x)$, displayed in the bottom panel of Fig. 2, is thus the key quantity to elucidate whether the strategy with no resetting in the bulk is optimal or not. In this sense, it plays a role similar to the function $M(x_{T})$, Eq. (5), shown in the top panel of Fig. 2, for the homogeneous resetting case—but with an additional dependence on the resetting position $x$. For those points $(x_{T},x)$ at which $M_{s}(x_{T},x)<0$, nonresetting in the bulk is not the best strategy for a target perfectly localized at $x_{T}$: introducing an infinitesimal delta kick at $x$ makes $\overline{\tau}$ decrease with respect its value with no resetting. Interestingly, there is an interval of “small” target positions, $0\leq x_{T}\leq x_{T}^{(0)}$, where $M_{s}<0$ for all $x$: this means that the optimal strategy for a target distribution with finite support in $(0,x_{T}^{(0)})$ involves nonvanishing resetting in the bulk regardless of further details of the target distribution. Conversely, there is not an interval of “large” target positions above which $M_{s}$ is positive for all $x$ and, thus, we cannot guarantee that a finite-support distribution leads to suppression of resetting in the bulk as the optimal strategy. However, looking at the figure, it is clear that the target positions that contribute to $r(x)=0$ as the optimal strategy are those close to the boundaries, i.e., $|x_{T}|$ close to unity—as already discussed in the simpler case of homogeneous resetting.
Now, we turn to studying the optimal resetting strategy for the $\beta$-family of target distributions (6). Since it is symmetric, $M_{s}(x_{T},x)$ is the key quantity. We write $\mu(x;\beta)$ because it depends on $\beta$—its analytical expression in terms of hypergeometric functions can be found in Sec. III of [40]. Nevertheless, we focus here on its sign, which is the main property for our purposes. For $\beta<1$, zero bulk resetting is optimal with $\mu(x;\beta)>0$ for $x\in(0,1)$. For $\beta>1$, $\mu(x;\beta)$ becomes negative close to the wall $x=1$, signaling that the no-resetting strategy is no longer the optimal one. Therefore, for the $\beta$-family (6), the flat distribution is the critical case, $\beta_{c}=1$; the critical $\beta$ is consistently smaller compared to the homogeneous case, $\beta_{c}\simeq 1.71$.
The optimal search strategy with bulk resetting cannot be derived analytically for any $P(x_{T})$, because Eq. (3) does not have an analytical solution for arbitrary $r(x)$. We therefore present numerical solutions, based on gradient descent algorithms [41], leading to the optimal $r^{*}(x)$—see Sec. IV of [40]. The obtained optimal strategy displays robust features for $\beta>1$, as shown in Fig. 3: there appears a value $x_{J}$ that separates zero resetting and nonzero resetting regions, i.e., $r^{*}(x)=0$ for $0<|x|<x_{J}$, whereas $r^{*}(x)\neq 0$ for $x_{J}<|x|<1$. Resetting is suppressed in $|x|<x_{J}$ to ensure uninterrupted exploration. For $|x|>x_{J}$, resetting is significantly inhomogeneous, displaying quite high values close to the walls. At $|x|=x_{J}$ the numerical results evidence the presence of a Dirac delta contribution to $r(x)$, which is corroborated by the use of different meshes. Interestingly, we report that the qualitative features of the profile for the optimal resetting obtained, i.e., the vanishing window, the peak and the high heterogeneity, are robust for other families with accumulation of probability in the center, as detailed in Sec. V of [40]. We further find that, for a specific family of target distributions, resetting strategies consisting in two symmetric delta peaks are the exact optimal solution, as detailed in Sec. V of [40].
Figure 3: Optimal heterogeneous resetting profiles in the bulk. Specifically, they have been obtained numerically for the target distributions of the family (6) corresponding to $\beta=0.5,2,3$. For $\beta\leq 1$, the optimal strategy consistently involves no resetting in the bulk. For $\beta>1$, for which nonzero bulk resetting is the optimal strategy, the features displayed—central region with no resetting, strongly heterogeneous behavior close to the wall—are robust. Symbols and lines respectively stand for meshes with $N=501$ and $N=2001$ nodes. The peaks become consistently higher and narrower when increasing the number of nodes in the mesh.
Conclusions.- In this Letter, we have shown that for one-dimensional search problems within a finite domain, resetting boundaries are always beneficial: if the searcher has not found the target upon reaching the wall, it is advantageous to be reset to the initial point. Then, we have analyzed the impact of the resetting strategy in the bulk, taking into account the role of the spatial distribution of the target. We have provided rigorous mathematical criteria to determine the optimality of the absence of resetting in the bulk, for both homogeneous and inhomogeneous resetting rates. These criteria, valid for arbitrary target distributions, have been further illustrated by considering a monoparametric family of such distributions—allowing us to shed further light on the physical reasons behind the optimality/nonoptimality of bulk resetting.
Our work paves new research lines in the study of search processes. On the one hand, we have shown that the spatial distribution of the target plays a key role in determining the optimal search strategy. This contrasts with the typical approach of optimizing the search strategy for a fixed, known position of the target: in an actual search process, the searcher does not know the position of the target. On the other hand, the optimal resetting strategy is highly nontrivial and inhomogeneous, even for the relatively “simple” one-dimensional case analyzed in this Letter. Trying to find an approximate expression for it is a promising research direction, as also is extending our analysis to searches in multidimensional domains.
Acknowledgments.- G. García-Valladares, C. A. Plata and A. Prados acknowledge financial support from Grant ProyExcel_00796 funded by Junta de Andalucía’s PAIDI 2020 programme. C. A. Plata and A. Prados acknowledge financial support from the applied research and innovation Project PPIT2024-31833, cofunded by EU–Ministerio de Hacienda y Función Pública–Fondos Europeos–Junta de Andalucía–Consejería de Universidad, Investigación e Innovación. A. Manacorda acknowledges financial support from the project “MOCA” funded by MUR PRIN2022 Grant No. 2022HNW5YL.
Data Availability Statement.- Numerical codes employed for generating the data and figures that support the findings of this study are openly available in the GitHub page of University of Sevilla’s FINE research group.